OpenStudy (anonymous):

Can someone Please Help? The side of a square is Three raised to the five halves power inches. Using the area formula A = s2, determine the area of the square.

OpenStudy (jchick):

(8^(3/2))^2 = 8^3 = 512

OpenStudy (mathmale):

"The side of a square is Three raised to the five halves power inches. Using the area formula A = s2, determine the area of the square." Let the side of a square be \[3^{\frac{ 5 }{ 2 }}.\]

OpenStudy (mathmale):

You want the area, and the formula for that is \[A=s^2\]

OpenStudy (mathmale):

Substitute the above expression for s into the formula for area and evaluate the area.

OpenStudy (jchick):

Thank you @mathmale I couldn't explain it very well.

OpenStudy (mathmale):

:)